How do i classify the linear system stated here : {x+4=y
{5y=5x+35
Can you help me by telling me what to do? I get lost .. I figured it out until the point of : 5y=(-5x-20)+35 ? I don't know the next step .. Help would be much appreciated .. Thanks :)

Answer by ptaylor(2198)

OK
x+4=y subtract 4 and also y from each side
x-y=-4----------------------eq1
5y=5x+35 subtract 5y and also 35 from each side
5x-5y=-35 divide each term by 5
x-y=-5-----------------------eq 2
Equations 1 and 2 represent parallel lines, therefore they are independent and inconsistent
Hope this helps---ptaylor